# Meta Engine: A Unified Semantic Query Engine on Heterogeneous LLM-Based Query Systems

### v1

@@ 5.3. Ablation Study
To understand the contribution of individual components in the Meta-Engine, we conduct a series of ablation experiments by selectively disabling specific modules. Each ablation isolates a design choice and reveals its functional role in the overall system. The results are shown in Table 3, 4 and 5.
5.3.1. Without Checker
In this setting, the complexity checker is disabled, and all queries are treated as complex by default. As a result, every input query is decomposed into sub-queries, regardless of whether decomposition is actually necessary.
@@ 5.3.1. Without Checker
Without checker, the result quality is usually lower in most cases. Although there are a few exceptions (Table 3), this does not affect the necessity of the checker, as checker improves the end-to-end efficiency significantly. As in Table 6, decomposing all queries without the checker could raise up to 70% more processing time. By preventing decomposing simple queries, a non-trivial portion of queries are processed in only one pass.
5.3.2. Without Decomposer
In the w/o Decomposer setting, the system never performs query decomposition. All original questions are considered non-complex and are passed directly into the pipeline as a single sub-query.
@@ 5.3.2. Without Decomposer
Removing decomposer makes the system overall perform worse, proving the effectiveness of query decomposition. There is only one exception in Table 4. This is because ManyModalQA has many single-hop queries, so decomposing everything may raises extra noise.
5.3.3. Without Aggregator
In this experiment, the final answer aggregator is disabled. While the system still decomposes the query and executes sub-queries iteratively, the final output is taken directly from the last sub-query’s answer, rather than being aggregating from all intermediate results.
@@ 5.3.3. Without Aggregator
In all cases the result quality drops significantly, especially F1-score score and Semantic Hit. This is because our aggregator comprehend the sub-queries results and ensure the final result is concise and more accurate. Hit score may be better without aggregator, but this is simply because the system tends to output long result, which has a larger chance to hit the ground-truth keywords but is actually not a correct answer.
5.3.4. Single Adapter (LOTUS)
We also evaluate the impact of the number of adapters. In the Single Adapter (LOTUS) configuration (the last row in the three tables), the system uses only the LOTUS adapter and disables the router as there is only one adapter to route. All other components from checker to aggregator are used as normal.
@@ 5.3.4. Single Adapter (LOTUS)
Comparing to the full versions of Meta Engine, single adapter system shows a lower performance, but it still outperforms the corresponding baseline method. Specifically, single LOTUS adapter Meta Engine achieves much higher F1-score (39.76 vs 9.47) and Semantic Hit (57.14 vs 53.90) on MultiModalQA, and all three metrics are significantly better than the LOTUS baseline on M2QA: F1-score 46.58 vs 5.43, Hit 69.06 vs 46.11, Semantic Hit 65.27 vs 38.72. These mean that (1) more adapters do increase the overall performance of Meta Engine, proving that Meta Engine effectively combines the specialities from different semantic systems; (2) The architecture and workflow design of Meta Engine effectively enhances the query processing quality of the underlying semantic systems.
Table 6. Average Query Runtime comparison (in seconds) including all ablations.
Method ManyModalQA MultiModalQA M2QA
Lotus 6.89 7.55 6.13
Single-Model 7.08 8.01 6.87
LlamaIndex 4.15 5.99 4.81
DSPy 7.37 9.03 11.63
@@
Full 8.15 11.91 19.89
w/o Aggregator 6.45 11.76 17.00
w/o Decomposer 5.76 7.33 8.30
w/o Checker 13.04 19.26 23.58
@@
Full 8.72 13.08 21.01
w/o Aggregator 7.53 12.26 20.99
w/o Decomposer 6.12 9.02 10.46
w/o Checker 14.53 18.74 25.81
@@ 5.4. Efficiency Analysis
We measure end-to-end efficiency on some datasets and report in Table 6. The result shows that Meta Engine achieves competitive runtime across most datasets. On ManyModalQA and MultiModalQA, most Meta Engine variants show a similar or faster running speed than baselines like DSPy and Single-Model, and remain close to lightweight baselines such as LlamaIndex. Meta Engine without checker, as we mentioned in Section 5.3, usually raises the latency, so it has the longest runtime. Although the full version of Meta Engine is slightly slower in some cases like the StatRouter version on MultiModalQA, the overhead remains moderate.
We will further improve the efficiency in our future work. Specifically, we observe that the latency is mainly raised in the operator ranking and generating stage, and we also observe this is depending on the dataset:
the operator generation attempts averagely 1.05 - 1.13 iterations per query on MultiModalQA, wihle it attempts 1.16 - 1.28 on ManyModalQA, and 1.52 - 1.81 on M2QA, which is positively correlated with the runtime on these datasets. This indicates the operator ranking is a key to guaranteeing the efficiency. So we will improve its accuracy to minimize the iterations in our future work. In addition, we plan to further improve efficiency through parallel sub-query processing: some sub-queries may have no dependency between each other, so they can be processed parallely instead of sequentially, which will effectively resolve the efficiency bottleneck. We will also apply more optimizations such as parallel execution of multiple queries and result caching for similar queries or sub-queries.
@@ 6. Conclusion
In this paper, we presented Meta Engine, a unified semantic query engine that coordinates multiple LLM-based query systems to support complex multimodal tasks. Instead of relying on a single monolithic model or a fixed execution pipeline, Meta Engine decomposes complex queries into modality-aligned sub-queries, routes each sub-query to the most suitable specialized system, and aggregates intermediate results to produce accurate and coherent final answers.
Extensive experimental results across diverse multimodal benchmarks demonstrate that Meta Engine achieves superior end-to-end answer quality in most settings, particularly on tasks requiring multi-step reasoning and multi-hop evidence integration.
Overall, this work presents a system that explicitly decomposes queries and coordinates multiple specialized systems. By enabling flexible integration and coordination of LLM-based specialized systems, Meta Engine provides a practical and extensible foundation for building high-performance multimodal semantic query applications.

### v2

@@ 5.3. Ablation Study
To understand the contribution of individual components in the Meta-Engine, we conduct a series of ablation experiments by selectively disabling specific modules. Each ablation isolates a design choice and reveals its functional role in the overall system. The results are shown in Table 4, 5 and 6.
@@
Without Checker: In this setting, the complexity checker is disabled, and all queries are treated as complex by default. As a result, every input query is decomposed into sub-queries, regardless of whether decomposition is actually necessary. Without checker, the result quality is usually lower in most cases. Although there are a few exceptions (Table 4), this does not affect the necessity of the checker, as checker improves the end-to-end efficiency significantly. As in Table 7, decomposing all queries without the checker could raise up to 70% more processing time. By preventing decomposing simple queries, a non-trivial portion of queries are processed in only one pass.
@@
Without Decomposer: In the w/o Decomposer setting, the system never performs query decomposition. All original questions are considered non-complex and are passed directly into the pipeline as a single sub-query. Removing decomposer makes the system overall perform worse, proving the effectiveness of query decomposition. There is only one exception in Table 5. This is because ManyModalQA has many single-hop queries, so decomposing everything may raises extra noise.
@@
Without Aggregator: In this experiment, the final answer aggregator is disabled. While the system still decomposes the query and executes sub-queries iteratively, the final output is taken directly from the last sub-query’s answer, rather than being aggregating from all intermediate results. In all cases the result quality drops significantly, especially F1-score score and Semantic Hit. This is because our aggregator comprehend the sub-queries results and ensure the final result is concise and more accurate. Hit score may be better without aggregator, but this is simply because the system tends to output long result, which has a larger chance to hit the ground-truth keywords but is actually not a correct answer.
@@
Single Adapter (LOTUS): We also evaluate the impact of the number of adapters. In the Single Adapter (LOTUS) configuration (the last row in the three tables), the system uses only the LOTUS adapter and disables the router as there is only one adapter to route. All other components from checker to aggregator are used as normal. Comparing to the full versions of Meta Engine, single adapter system shows a lower performance, but it still outperforms the corresponding baseline method. Specifically, single LOTUS adapter Meta Engine achieves much higher F1-score (39.76 vs 9.47) and Semantic Hit (57.14 vs 53.90) on MultiModalQA, and all three metrics are significantly better than the LOTUS baseline on M2QA: F1-score 46.58 vs 5.43, Hit 69.06 vs 46.11, Semantic Hit 65.27 vs 38.72. These mean that (1) more adapters do increase the overall performance of Meta Engine, proving that Meta Engine effectively combines the specialities from different semantic systems; (2) The architecture and workflow design of Meta Engine effectively enhances the query processing quality of the underlying semantic systems.
Table 7. Average Query Runtime comparison (in seconds)
Method ManyModalQA MultiModalQA M2QA Text2Vis
Lotus 6.63 7.34 5.75 10.72
Single-Model 6.96 7.76 6.61 9.62
LlamaIndex 3.98 5.72 4.46 10.68
DSPy 7.13 8.68 11.15 13.27
@@
Full 7.76 10.59 15.91 18.06
W/O Checker 10.76 14.54 18.29 21.80
@@
Full 8.28 10.75 20.81 22.66
W/O Checker 11.67 16.11 23.70 24.86
@@ 5.4. Efficiency Analysis
We measure end-to-end efficiency on several datasets and report the results in Table 7. Particularly, we include the results of Meta Engine without complexity checker, to show the importance of dynamic query decomposition to the overall efficiency. On ManyModalQA and MultiModalQA, the full version of Meta Engine achieves competitive runtime to several baselines. For example, Meta Engine (MLRouter) runs in 7.76s on ManyModalQA, close to DSPy (7.13s) and Single-Model (6.96s); on MultiModalQA, Meta Engine (MLRouter) takes 10.59s, competitive to DSPy at 8.68s. On M2QA and Text2Vis, Meta Engine incurs higher latency (which is still acceptable like 15.91 VS 11.15 of DSPy), primarily due to the increased number of operator generation iterations required by the more complex queries in these datasets. Specifically, the operator generation stage averages 1.05–1.13 iterations per query on MultiModalQA, while it averages 1.16–1.28 on ManyModalQA, and 1.52–1.81 on M2QA, showing a clear positive correlation between iteration count and runtime. This suggests that improving the accuracy of the operator ranking mechanism is a promising direction for reducing latency.
As discussed in Section 5.3, removing the checker consistently increases runtime across all datasets (e.g., from 7.76s to 10.76s on ManyModalQA with MLRouter, and from 10.59s to 14.54s on MultiModalQA), as the maximum number of subqueries have to be executed for each query. This further confirms the importance and effectiveness of the dynamic query decomposition in both accuracy and efficiency. Additionally, the pipeline with statistic router is slower than that with learned router. This shows another advantage of learned router: it has a more accurate favor for the efficient adapter while pursuing the best quality.
@@ 6. Conclusion
In this paper, we presented Meta Engine, a unified framework for building semantic query pipelines that coordinates multiple LLM-based query systems to support complex multimodal tasks. Meta Engine provides easy-to-use APIs and flexible and composable components (from query decomposer to result aggregator) to minimize the users’ cost to customize their pipelines and integrate multiple LLM systems. Extensive experimental results across diverse multimodal datasets demonstrate that the pipelines built on Meta Engine achieve superior query processing quality while maintaining competitive efficiency comparing to state-of-the-art LLM-based query systems.
7. Self-Declaration of Related Submissions
We confirm that there are no related submissions by the same authors currently under review at PVLDB or any other venue.